An athlete participates in a discus throw competition during an Olympiad that takes place in a city where acceleration due to gravity is 9.7 m/s^2. He throws the discus with an initial velocity of 20 m/s at an angle of 60 degrees from the vertical. Neglecting air resistance and the height of the discus at the point of release, what is the range of the discus?

PLEASE HELP. I really need a GOOD explanation on the concept and steps. Pleaseeeeee!!
Answer #1

Range R = u^2 Sin2@/g

Here U = 20m/s

@= 60°

R = 400×Sin2(60) / 9.7 = 35.7m

=) R = 35.7m